OpenStudy (christos):

modulos math, Which of the following gives the 3rd root of 92 modulo 187? (Note that 187=11⋅17.)

OpenStudy (anonymous):

okay

hartnn (hartnn):

3rd root ? or \(92^3\) ?

OpenStudy (anonymous):

not making sense

OpenStudy (dumbcow):

92 mod 187 = 92 , the remainder is simply 92 in this case

hartnn (hartnn):

i am sure, she knows that... lets assume its \(\Large 92^3 mod ~187\)

OpenStudy (christos):

ok

hartnn (hartnn):

can you post the screenshot of the question ?

OpenStudy (christos):

ok hold on one sec

OpenStudy (christos):

what is the question telling me the prime factors for 187 ( 11 * 17) ? and what does it mean by 3rd root ?

OpenStudy (christos):

is it like x^3 = 92 modulo 187? and I have to find x ?

hartnn (hartnn):

I am not sure, i'll wait for other experts to look into it...sorry

OpenStudy (christos):

its ok np

OpenStudy (dumbcow):

in that case i think you would just take 3rd root of 92

OpenStudy (loser66):

I would like to see "the following", please.

hartnn (hartnn):

i thought mod arithmetic works with integers only

OpenStudy (anonymous):

sumbody temme what she wants !!

OpenStudy (christos):

should very well be the one hartnn assumed , but I am more like looking for the reasoning

hartnn (hartnn):

all i can say is : its not 2nd option :P

OpenStudy (christos):

xD

hartnn (hartnn):

is it asking for 3rd primitive root ... :O

OpenStudy (anonymous):

hartnn seems to know

OpenStudy (christos):

most likely

OpenStudy (christos):

but wats dat

OpenStudy (christos):

I need to use the prime factors of 187 somehow

hartnn (hartnn):

@ganeshie8 @mathmath333

ganeshie8 (ganeshie8):

you want to solve \[x^3 \equiv 92 \pmod {187}\]

hartnn (hartnn):

what does it mean by 'give 3rd root of 92 modulo 187' ?

ganeshie8 (ganeshie8):

find some integer whose cube is 92

OpenStudy (christos):

it doesnt exist

ganeshie8 (ganeshie8):

it might exist in modulo 187

ganeshie8 (ganeshie8):

looking at your options, it does exists

OpenStudy (christos):

how can you pin point so fast though

OpenStudy (christos):

looking at the options i mean

hartnn (hartnn):

what we use (decimal) is modulo 10 ?

ganeshie8 (ganeshie8):

we need to work the root was just saying you dont have an option like "root does not exist"

OpenStudy (christos):

I don't @hartnn im new to modular arithmetic , might very well be

ganeshie8 (ganeshie8):

lets use the hint maube..

OpenStudy (christos):

hmm

OpenStudy (christos):

by the way right answer is: https://www.dropbox.com/s/fx0z4egq4ket670/Screenshot%202014-12-30%2021.22.15.jpg?dl=0

OpenStudy (anonymous):

really got to brush up my modulos

OpenStudy (christos):

just in case it helps with anything regarding the process

ganeshie8 (ganeshie8):

thats the right answer but idk any easy method to explain you

OpenStudy (christos):

got it, thanks a lot

ganeshie8 (ganeshie8):

you will need to know euler phi function and few other things to make sense of above ^ let me see if i can figure out a solution using basic mod properties

OpenStudy (christos):

nah I know phi function and the others, I understand perfectly now

ganeshie8 (ganeshie8):

fixed typo : \[\phi(187) = \phi(11\cdot 17) = 10\cdot 16 = 160\] so we have \[92^{160} \equiv \color{Red}{1} \pmod {187}\] \[x^3\equiv 92\times 92^{2\times 160} \pmod {187}\] \[x^3\equiv 92^{2\times 321} \pmod {187}\] \[x\equiv 92^{ 321/3} \pmod {187}\]

OpenStudy (christos):

thanks a bunch

ganeshie8 (ganeshie8):

you're good then :)

hartnn (hartnn):

*sigh* some day :P

OpenStudy (christos):

xD

ganeshie8 (ganeshie8):

phi function is just an extension to fermat little theorem

ganeshie8 (ganeshie8):

Fermat little theorem : whenp is prime and gcd(a,p)=1 below holds \[a^{p-1}\equiv 1 \pmod {p}\]

ganeshie8 (ganeshie8):

euler generalization of fermat little theorem using phi function : \[a^{\phi(n)}\equiv 1 \pmod {n}\] this is true for all integers n, it need not be prime anymore however gcd(a,n)=1 is still required

ganeshie8 (ganeshie8):

\(\phi(n)\) gives the number of integers less than \(n\) that are coprime to \(n\)

hartnn (hartnn):

\(\phi (n) =n-1\) is tru only if n is prime

ganeshie8 (ganeshie8):

yes when n is prime, euler generalization becomes fermat little thm

hartnn (hartnn):

ok thanks :) i'll dig more :)